OpenStudy (gabby48):

What would you do to isolate the variable x+5=-12

OpenStudy (jackthegreatest):

how do you move the 5 from left to the right side?

OpenStudy (jackthegreatest):

@Gabby48 any ideas?

OpenStudy (jackthegreatest):

\[x + 5 -5 = -12 -5\]

OpenStudy (jackthegreatest):

does that help @Gabby48

OpenStudy (quickstudent):

It's very easy :) Here is how you isolate x: x + 5 = -12 x + 5 - 5 = -12 - 5 x = -17

OpenStudy (nduong21):

1st step you need to isolate the variable\[x +5 = -12\] subtract 5 from both sides\[x=-12 -5\] do the math \[x=-17\]